OpenStudy (anonymous):

A cylinder, with base of raidus x cm and height h cm, has the same volume as a cuboid with length x cm, width x cm and height (x + 2) cm. Find an expression for h in terms of x and pi, simplifying your answer.

OpenStudy (anonymous):

V of Cylinder = pi(r^2)h V of cuboid = l*w*h plug in the terms from the question. &make the volumes equal to each other.

OpenStudy (anonymous):

pi(x^2)h = x(x)(x+2) h = x(x)(x+2) / pi(x^2) simplify.